OpenStudy (anonymous):

slope:-1/9 ordered pair (0,3)

OpenStudy (wolf1728):

We start with the formula b = y - mx b = 3 -(1/9)•0 b = 3 y = mx + b y = -(1/9) + 3

OpenStudy (wolf1728):

CORRECTION y = -(1/9)x + 3